OpenStudy (anonymous):

Let g(x)=2x and h(x)=x^2 +4 (g of h)(0)

OpenStudy (anonymous):

\[g(h(0)) = g(0^{2} +4)\]

OpenStudy (anonymous):

= g(4)

OpenStudy (anonymous):

g(4) = 2(4)= 8